# Leveraging Generative AI for Smart City Digital Twins: A Survey on the Autonomous Generation of Data, Scenarios, 3D City Models, and Urban Designs

### v1

@@ Abstract
The digital transformation of modern cities by integrating advanced information, communication, and computing technologies has marked the epoch of data-driven smart city applications for efficient and sustainable urban management. Despite their effectiveness, these applications often rely on massive amounts of high-dimensional and multi-domain data for monitoring and characterizing different urban sub-systems, presenting challenges in application areas that are limited by data quality and availability, as well as costly efforts for generating urban scenarios and design alternatives. As an emerging research area in deep learning, Generative Artificial Intelligence (AI) models have demonstrated their unique values in data and code generation. This survey paper aims to explore the innovative integration of generative AI techniques and urban digital twins to address challenges in the realm of smart cities in various urban sectors, such as transportation and mobility management, energy system operations, building and infrastructure management, and urban design. The survey starts with the introduction of popular generative AI models with their application areas, followed by a structured review of the existing urban science applications that leverage the autonomous capability of the generative AI techniques to facilitate (a) data augmentation for promoting urban monitoring and predictive analytics, (b) synthetic data and scenario generation, (c) automated 3D city modeling, and (d) generative urban design and optimization. Based on the review, this survey discusses potential opportunities and technical strategies that integrate generative AI models into the next-generation urban digital twins for more reliable, scalable, and automated management of smart cities.
@@
Generative artificial intelligence , smart city , urban digital twin , deep learning
@@ 1 Introduction
The digital transformation of cities, propelled by advancements in the Internet of Things (IoT), Information and Communication Technologies (ICT), and 5G cellular networks, is leading to the generation of substantial amounts of urban data (Allam and Dhunny, 2019; Javed et al., 2022), enabling the monitoring and situational awareness of various urban sub-systems (Hossain et al., 2018). Driven by the big urban data, smart city applications have reshaped the management of the environment and urbanization in different urban sectors to proliferate the sustainability, productivity, livability, and economic growth in cities across the world (Ullah et al., 2020; Kim et al., 2020a). Examples of these urban sectors include intelligent transportation (Xiong et al., 2012; Zhao and Jia, 2021), urban mobility (Xu et al., 2023b; Paiva et al., 2021), energy smart buildings (Khajenasiri et al., 2017; Francisco et al., 2020), water management and distribution (Karwot et al., 2016), and power grid operations (Omitaomu and Niu, 2021). With the recent advent of Artificial Intelligence (AI) and Digital Twin (DT) technologies, traditional machine learning and deep learning techniques have been ubiquitously applied in smart city applications to support advanced urban monitoring and situational awareness, facilitate complex decision-making, as well as the automation and autonomy in complex urban management processes. (Allam and Dhunny, 2019; Luckey et al., 2021; Cugurullo, 2020; Farsi et al., 2020). Many smart city applications adopt the paradigm of emerging digital twins to develop urban digital twins (i.e., smart city digital twins or digital twin cities) (Shahat et al., 2021; Dembski et al., 2020). The original concept of “Digital Twin” focuses on the development of virtual models of real-world objects and processes in a computing environment to support advanced manufacture and analytics (Ivanov et al., 2020), and was applied to the urban science sector to create urban digital twins, which are digital replicas of cities and their sub-urban system for intelligent urban management and planning, as well as for enhancing smart city servers. These urban digital twins leverage advanced computing and cyberinfrastructure technologies to promote and automate smart city services (Shirowzhan et al., 2020; Xu et al., 2023b). Urban digital twins represent a groundbreaking leap in urban science and planning, especially in the realms of mobility management, building energy optimization, and urban design. By creating detailed, dynamic digital replicas of urban environments, these tools enable planners and scientists to simulate and analyze complex systems with unprecedented accuracy and efficiency in many urban sectors, such as transportation and mobility management (Xu et al., 2023b), building energy systems (Bortolini et al., 2022), smart grids (Cioara et al., 2021), and infrastructure management (Callcut et al., 2021; Sabri et al., 2023). In the broader context of urban design, digital twins offer a holistic view of the urban ecosystem, enabling planners to create more livable, resilient, and adaptive urban spaces (Dembski et al., 2020). The integration of urban digital twins in these sectors not only streamlines planning processes but also fosters examining scenarios and innovative solutions to contemporary urban challenges, paving the way for smarter, more sustainable cities. Some examples are using urban digital twins for water-sensitive urban design (Langenheim et al., 2022), and development envelop control analysis (Sabri et al., 2022).
@@ 1.1 Challenges in Urban Digital Twins
Although there have been significant advancements in AI, communication, and computing technologies, urban (or smart city) digital twins that rely on big urban data still encounter various methodological and technical challenges in many aspects:
Data Availability and Quality:
@@
Hypothetical Urban Scenarios:
@@
There is an increasing trend for smart city initiatives to develop 3D city models as critical components of urban digital twins to support data integration and visualization, interactive planning and simulation, real-time monitoring and response, and public engagement (Deng et al., 2021; Farsi et al., 2020). However, developing a comprehensive 3D city model to digitize a large urban area in a computing system can be a costly, repetitive, and labor-intensive endeavor (Botín-Sanabria et al., 2022; Singh et al., 2021), which involves massive modeling and gamification of urban entities (e.g., buildings and infrastructure) as 3D models with different levels of details. Additionally, constructing 3D terrain, underground urban objects, and vegetation using geomatics and surveying data can also be a time-consuming effort (Xu et al., 2023a).
Urban Design:
@@
Built upon the previous data analytics, simulations, and digital twin applications, the planning and management of smart cities often involve top-down and multi-scale urban design processes, which are responsible for the configuration of land-use and land-cover, transportation network, buildings, and infrastructure (Van Nes and Yamu, 2021; Pérez-Martínez et al., 2023). These processes are often repetitive, expensive, and time-consuming, involving multiple design alternatives and complex decision-making for design optimization. Therefore, it is challenging to scale up the design processes to large urban areas, which often have distinct design requirements (e.g., green space and accessibility).
@@
Given the complexities, crafting wide-ranging and insightful solutions to urban issues from an all-encompassing viewpoint is proving to be progressively more expensive and demanding in terms of effort. These solutions require the analysis of complex urban systems — including transportation, energy, building, infrastructure, and water — as interconnected subsystems at the ground, above-ground, and underground levels. The selection of these urban subsystems is based on previous literature and urban digital applications (Ferré-Bigorra et al., 2022, 2022; Hämäläinen, 2021). Consequently, there is a growing need for developing scalable and autonomous solutions. Such solutions would employ generative AI techniques to enhance the availability and quality of urban data, automate the generation of various scenarios, support the development of 3D city models, and automate complex urban design and optimization processes (Wang et al., 2023a; Kirwan and Zhiyong, 2020).
Figure 1: A summary of generative AI-related publications in major urban management sector using Scopus database (Search conducted on Jan 10th 2024)
@@ 1.2 Motivation
In the realm of deep learning, Generative Artificial Intelligence (AI) has emerged as a groundbreaking field, showcasing its unique ability to create novel content from realistic data. This capability is increasingly relevant for enhancing smart city applications. Despite the rapid proliferation of generative AI in various urban science and management sectors, there remains a notable absence of a thorough review encompassing mainstreaming generative AI techniques and their contributions to urban and environmental management within the context of developing smart cities through digital twin models. This survey paper aims to fill this gap by providing a structured overview of current generative AI models and their applications in diverse urban management sectors, particularly those not yet explored from the smart city digital twin perspective. Drawing upon recent literature and proposed agendas, the focus of this survey is primarily on key smart city application areas, such as transportation and mobility management, energy system operations, urban water management, and building and infrastructure management (Shahat et al., 2021; Dembski et al., 2020). This survey searched and discovered the publication history and trend of generative AI-powered studies in these application areas using the Scopus database, and present our discovery in Figure 1. In addition, this paper reviews the capability of generative AI techniques in urban analysis (e.g., population and morphology), 3D city modeling, and generative urban design. Organized around the limitations of urban digital twins in data availability and quality, scenario generation, 3D city modeling, and urban design, this survey critically examines the potential integration of generative AI models to enhance the capability, reliability, and efficiency of urban digital twins. Finally, the paper discusses future challenges and opportunities from a software engineering perspective, offering insights into the evolving landscape of smart city development.

### v2

# Leveraging Generative AI for Urban Digital Twins: A Scoping Review on the Autonomous Generation of Urban Data, Scenarios, Designs, and 3D City Models for Smart City Advancement

@@ Abstract
The digital transformation of modern cities by integrating advanced information, communication, and computing technologies has marked the epoch of data-driven smart city applications for efficient and sustainable urban management. Despite their effectiveness, these applications often rely on massive amounts of high-dimensional and multi-domain data for monitoring and characterizing different urban sub-systems, presenting challenges in application areas that are limited by data quality and availability, as well as costly efforts for generating urban scenarios and design alternatives. As an emerging research area in deep learning, Generative Artificial Intelligence (GenAI) models have demonstrated their unique values in content generation. This paper aims to explore the innovative integration of GenAI techniques and urban digital twins to address challenges in the planning and management of built environments with focuses on various urban sub-systems, such as transportation, energy, water, and building and infrastructure. The survey starts with the introduction of cutting-edge generative AI models, such as the Generative Adversarial Networks (GAN), Variational Autoencoders (VAEs), Generative Pre-trained Transformer (GPT), followed by a scoping review of the existing urban science applications that leverage the intelligent and autonomous capability of these techniques to facilitate the research, operations, and management of critical urban subsystems, as well as the holistic planning and design of the built environment. Based on the review, we discuss potential opportunities and technical strategies that integrate GenAI models into the next-generation urban digital twins for more intelligent, scalable, and automated smart city development and management.
@@
Generative artificial intelligence , smart city , urban digital twin , 3D City Modeling , urban planning , deep learning
@@ 1 Introduction
Planning and managing urban environments often require complex, interdisciplinary efforts that encompass various urban subsystems, including transportation and mobility, urban energy, water supply and distribution, as well as buildings and infrastructure (Fedorowicz et al., 2020). With the recent advent of the Internet of Things (IoT), Information and Communication Technologies (ICT), and 5G cellular networks, cities worldwide have been digitally transformed to generate substantial amounts of urban data, opportunities for more intelligent and efficient smart city applications to enhance urban sustainability, productivity, and livability (Hossain et al., 2018). This transformation also has reshaped the management of the various urban subsystems through the integration of Artificial Intelligence (AI) and emerging digital twin technologies, which, in combination, created a new paradigm for developing urban digital twins, which is also known as digital twin cities, city digital twins, and smart city digital twins, to promote smart city services (Shahat et al., 2021; Dembski et al., 2020).
The original digital twin concept focuses on the development of virtual models of real-world objects and processes in a computing environment to support advanced manufacture and analytics (Ivanov et al., 2020), and was initially proposed to support advanced manufacturing. Similarly, the urban digital twin paradigm replicates different urban subsystems in the digital environment to enable a variety of advanced capabilities, including real-time situational awareness and predictive analytics of complex urban dynamics, as well as the informed decision support, optimization, and automation of sophisticated urban management tasks through the integration of advanced visualization and visual analytics (Beckett, 2022; Xu et al., 2022a), and cyber-physical integration (Allam and Dhunny, 2019; Luckey et al., 2021), and City Information Models (CIMs) (Shi et al., 2023).
1.1 Digital Twins for Smart Cities
Following this paradigm, numerous operational urban digital twins have been developed and implemented to support the management of various real-world urban subsystems, including urban mobility (Xu et al., 2023b), building operations (Bortolini et al., 2022), urban energy and power grids (Cioara et al., 2021; Jafari et al., 2023), and infrastructure (Callcut et al., 2021; Sabri et al., 2023). In the broader context of urban planning and design, digital twins offer a holistic view of the urban ecosystem, enabling planners to create more livable, resilient, and adaptive urban spaces (Dembski et al., 2020). Examples include the use of urban digital twins for water-sensitive urban design (Langenheim et al., 2022) and generative development envelope control analysis (Sabri et al., 2022). The incorporation of urban digital twins in these sectors not only streamlines planning processes but also facilitates the examination of scenarios and the development of innovative solutions to contemporary urban challenges, paving the way for smarter, more sustainable, human-centered, and participatory cities.
By examining contemporary digital twin studies, We have created Table 1 to provide an overview of previous successful urban digital twin applications developed in various urban sectors. The table utilizes major application areas defined by recent review articles on digital twins for various urban subsystems (Ivanov et al., 2020), including transportation (Bao et al., 2021; Wang et al., 2022b), urban energy (Jafari et al., 2023; Pan et al., 2020), urban water (Ramos et al., 2023b), and building and infrastructure (Al-Sehrawy et al., 2021). We also explored the utilization of digital twin on the planning, design, and management of the urban environment as the system of multiple subsystems from a holistic aspect (Al-Sehrawy et al., 2021).
Table 1: An overview of urban digital twin examples in various urban application areas.
Application Area Examples
Transportation and Mobility Traffic Simulation (Kušić et al., 2023; Saroj et al., 2021; Xiong et al., 2022) Traffic Control Optimization (Xu et al., 2023b; Kamal et al., 2024; Kumarasamy et al., 2024) Traffic Data Prediction (Hu et al., 2021; Nie et al., 2023; Kumar et al., 2018) Connected Automated Vehicles (Fan et al., 2022a; Du et al., 2021; Ali et al., 2023; Wágner et al., 2023) Traffic Safety (Irfan et al., 2024; Fu et al., 2024; Lv et al., 2022b; Wang et al., 2022a; Lv et al., 2022a) Transportation Asset Management (Gao et al., 2021; Wu et al., 2022b; Adibfar and Costin, 2022) Advanced Driver-Assistance Systems (Wang et al., 2022b, 2020b; Schwarz and Wang, 2022) Parking Management (Chomiak-Orsa et al., 2023; Zou et al., 2023; Liu et al., 2023; Shao et al., 2022)
Urban Energy Grid Operation (Jafari et al., 2023; Seong Lok et al., 2024; Sifat et al., 2023; Nordzi et al., 2022) Energy Management (Srinivasan et al., 2020; Xu and Liu, 2024; Francisco et al., 2020) Anomaly and Fault Detection (Danilczyk et al., 2021; Jain et al., 2019; Joseph et al., 2018) Renewable Energy (Xu and Gong, 2023; Belik and Rubanenko, 2023; Agostinelli et al., 2021)
Urban Water Smart Water Grid (Wu et al., 2023; Ramos et al., 2023a; Giudicianni et al., 2020) Water Supply and Distribution (Conejos Fuertes et al., 2020; Karmous-Edwards et al., 2019; Martínez Alzamora et al., 2021) Drainage and Stormwater System (Bartos and Kerkez, 2021; Sharifi et al., 2024) Wastewater Management (Hallaji et al., 2021; Wang et al., 2024; Jiang et al., 2022)
Building and Infrastructure Building Operation (Zhao et al., 2022; Hodavand et al., 2023; El Mokhtari et al., 2022) Building Information Modeling (Qiuchen Lu et al., 2019; Sepasgozar et al., 2020; Sacks et al., 2020) City Information Modeling (Shi et al., 2023; Shariatpour et al., 2024; Petrova-Antonova and Ilieva, 2021) Infrastructure Management (Al-Sehrawy et al., 2021; Broo et al., 2022; Shirowzhan et al., 2020; Abdeen et al., 2023; Yu and He, 2022)
Urban Planning and Design (Yang and Kim, 2021; Schrotter and Hürzeler, 2020; White et al., 2021; Kumalasari et al., 2023; Beckett, 2022; Xu et al., 2023a)
@@ 1.2 Challenges in Urban Digital Twins
Although there have been significant advancements in AI, communication, and computing technologies, urban digital twins, driven by the big data, still encounter various methodological and technical challenges in many aspects. We have outlined 6 major challenges based on the urban digital twin articles reviewed in the previous subsection. These challenges are labeled from C1 to C5 in the following list.
C1. Data Availability and Quality:
@@
C2. Hypothetical Urban Scenarios:
@@
There is an increasing trend for smart city initiatives to develop 3D city models as critical components of urban digital twins to support data integration and visualization, interactive planning and simulation, real-time monitoring and response, and public engagement (Deng et al., 2021; Biljecki et al., 2015). However, developing a comprehensive 3D city model for large urban areas is a costly, repetitive, and labor-intensive endeavor (Botín-Sanabria et al., 2022; Singh et al., 2021), which involves clarification of needed urban entities (e.g., buildings, transportation infrastructure, green areas, water bodies) the levels of details, the geometric accuracy, the type representations (vector or voxel), the validity specifications (e.g. volmetric objects have to be water-tight), the involvement of photo-realistic texture (as in reality models), the data structure for management of the 3D models as well as the allowed data types (e.g. surfaces, volumes, B-splines, parametric shapes). Additionally, constructing valid 3D model, preventing intersections and discrepancies between terrain, underground and above ground urban objects, can also be a time-consuming effort (Xu et al., 2023a). The integration of data for 3D city models relies on clear specifications of vocabularies, semantic definitions, explicit relationships and constrains. Despite the international standardization efforts and developed transformation and integration tools, many aspects of creating 3D City models require manual or semi-automatic cleaning, adjustment and conversion.
C4. Urban Design Scalability:
@@
Built upon the previous data analytics, simulations, and digital twin applications, the planning and management of smart cities often involve top-down and multi-scale urban design processes, which are responsible for the configuration of land-use and land-cover, transportation network, buildings, and infrastructure (Van Nes and Yamu, 2021; Pérez-Martínez et al., 2023). These processes are often repetitive, expensive, and time-consuming, involving multiple design alternatives and complex decision-making for design optimization. Therefore, it is challenging to scale up the design processes to large urban areas, which often have distinct design requirements (e.g., green space and accessibility). Subsequently, there is a need to developed automated generative co-design tools that can automated repeated and laborious processes.
C5. Data Sovereignty Requirements:
Data ownership, data colonialism, and data residency are key concerns related to data sovereignty. These issues can significantly constrain and restrict the sharing and usage of urban data in smart city applications (Rijshouwer et al., 2022; Rogan et al., 2019). Data ownership complexities arise when multiple stakeholders, such as governments, private companies, and individuals, claim rights over the data, leading to disputes and hindering collaborative efforts (Löfgren and Webster, 2020). Data colonialism exacerbates these challenges by creating power imbalances where external entities exploit local data resources without equitable benefits to the local communities, fostering mistrust and resistance to data sharing (Leese et al., 2022). Data residency requirements further complicate matters, as strict regulations mandate that data must be stored and processed within specific geographic boundaries, limiting cross-border data flows and integration (Navarrete et al., 2009). These issues collectively impede the seamless exchange of urban data, essential for developing and deploying innovative smart city solutions, ultimately restricting the potential for enhanced urban management and improved quality of life for residents.
C6. Public Participation and Community Engagement
: Engaging citizens and residents without an urban science background in participatory and human-centered urban planning and smart city management presents a significant challenge (Senior et al., 2023). The complexity of urban planning concepts and the technical details of smart city initiatives can be daunting for individuals lacking specialized knowledge and access to specialized software tools, creating a technical barrier to effective engagement (Chourabi et al., 2012; Tan and Taeihagh, 2020). Communication barriers further complicate this issue, as translating technical jargon into accessible and meaningful information is challenging. In this regard, it is essential to build a human-AI partnership through AI agents or chatbots that possess advanced human language understanding, domain knowledge, and the ability to autonomously perform specialized tasks such as knowledge base querying, data analysis, and simulation. These capabilities can be enabled through APIs integrated with urban digital twins. The AI agent should be able to directly communicate with various stakeholders, understand their individual needs, and formulate citizen-centric planning and management strategies. This approach ensures that the AI agents can facilitate meaningful engagement, making urban planning more inclusive and responsive to the community’s diverse needs.
@@
Given the complexities, crafting wide-ranging and insightful solutions to urban issues from an all-encompassing viewpoint is proving to be progressively more expensive and demanding in terms of effort. These solutions require the analysis of complex urban systems — including transportation, energy, building, infrastructure, and water — as interconnected subsystems at the ground, above-ground, and underground levels (Gallotti et al., 2021). The selection of urban subsystems discussed in this paper is guided by previous literature and urban digital twin applications (Ferré-Bigorra et al., 2022; Hämäläinen, 2021), with the topic further refined by examining successful urban digital twin implementations detained in Section 1.1. Consequently, there is a growing need for developing scalable, autonomous, and intelligent solutions. Such solutions would employ GenAI techniques to enhance the availability and quality of urban data, automate the generation of various scenarios, support the development of 3D city models, and automate complex urban design and optimization processes (Wang et al., 2023b; Kirwan and Zhiyong, 2020). Some GenAI models excel in understanding human languages and handling text, providing opportunities to create responsive AI agents that facilitate communication and co-design in planning and management. These AI agents enhance community participation and engagement by interpreting and generating natural language, making it easier for stakeholders to contribute their ideas and feedback in the planning process (Fares, 2023). This leads to more inclusive, transparent, and collaborative decision-making, ultimately resulting in urban designs that better reflect the needs and preferences of the community.
@@ 1.3 Motivation
In the realm of deep learning, Generative Artificial Intelligence (GenAI) has emerged as a groundbreaking field, showcasing its unique ability to create novel content from realistic data. This capability is increasingly relevant for enhancing smart city applications. Despite the rapid proliferation of GenAI in various urban science and management sectors, there remains a notable absence of a thorough review encompassing mainstreaming GenAI techniques and their contributions to urban and environmental management within the context of developing smart cities through digital twin models.
This survey paper aims to fill this gap by providing a structured overview of current GenAI models and their applications in diverse urban management sectors, particularly those not yet explored from the urban digital twin perspective. Drawing upon recent literature and proposed agendas, the focus of this survey is primarily on key smart city application areas that include transportation, energy system, urban water, and building and infrastructure (Shahat et al., 2021; Dembski et al., 2020).
1.4 Review Method
Given the interdisciplinary nature of this paper’s objectives, which encompass diverse topics from AI, urban informatics, and the built environment, we employed the scoping review methodology (Mak and Thomas, 2022) to systematically map emerging concepts and identify research gaps and applications related to the synergy between GenAI models and urban digital twin applications in the smart city sector. This methodology provides a knowledge synthesis that is particularly useful for several specific scenarios involving emerging research areas, interdisciplinary topics, broad research questions, and complex or heterogeneous literature (Levac et al., 2010). To formulate this survey and identify relevant literature, we follow the scoping review methodological framework outlined by Arksey and O’Malley (2005) through a series of procedures described in the following subsections.
1.4.1 Preliminary Search
It is important to define clear and broad research questions that will guide the selection of the domain, scope, and focus areas of the literature review. Before formulating these questions, Mak and Thomas (2022) suggests conducting a preliminary search of the literature to assess the feasibility of the review with regard to the following aspects:
  1. 1.
    The breadth of the research question.
  2. 2.
    Whether a scoping review on the topic has already been conducted.
  3. 3.
    If there is sufficient literature to warrant a scoping review.
Based on these aspects, we conducted a preliminary search. Based on our topic and research motivation described in Section 1.3, the breadth of our research scope encompasses a variety of domains and disciplines. The concept and applications of urban digital twins, also known as city digital twins, pertain to the fields of urban informatics, urban computing, city information modeling (CIM), and computational urban science (Shi et al., 2022a; Deng et al., 2021). Meanwhile, GenAI algorithms and applications represent an emerging area within the field of deep learning.
To determining whether a scoping review on this topic has already been conducted, we created a list of predefined syntax terms and conducted a literature search across major academic databases, including Google Scholar, Scopus, and IEEE Xplore. Our predefined search terms included ‘Generative Artificial Intelligence’, ‘Generative AI’, ‘GAI’, ‘GenAI’, ‘Generative’, ‘Generative Deep Learning’, ‘Digital Twin Cities’, and ‘City Digital Twins’. The search was performed on paper topics, including article titles, abstracts, and keywords. This search did not identify any previous review or survey articles on the subject.
To verify the quantity and availability of relevant literature, we developed a set of predefined search syntaxes tailored to the parent disciplines of ’Digital Twin Cities’ and their synonyms. These parent disciplines include ’Smart Cities,’ ’City Information Modeling,’ ’Urban Computing,’ and ’Urban Informatics,’ as identified by previous studies (Zheng et al., 2014; Xia et al., 2022). We combined these predefined syntaxes with GenAI search terms to conduct preliminary searches on the Scopus and IEEE Xplore academic databases. The Scopus search returned over 150 articles, while IEEE Xplore yielded 400 articles. Our preliminary searches confirmed that there is sufficient literature to justify proceeding with the proposed scoping review. For the subsequent in-depth literature review on individual GenAI applications, we conducted multiple searches across the two databases using refined search syntax that specifies not only the generic GenAI and smart city concepts but also detailed types of GenAI models and specialized urban research and management topics.
1.4.2 Identifying the Research Questions
Based on the results of the preliminary search, we have formulated 4 research questions and conducted the necessary preliminary searches to seek their answers. The answers to these fundamental research questions will guide the direction and structure of this scoping paper in a step-by-step manner.
Q1: How urban digital twins are applied to facilitate smart city advancement?
Based on previous reviews in urban science and urban planning, urban digital twins are developed to enhance the operation, management, and research of various urban subsystems, including transportation, energy, water, buildings, and infrastructure (Lu et al., 2021; Pesantez et al., 2022; Broo et al., 2022). They also play a crucial role in supporting critical urban planning tasks (Schrotter and Hürzeler, 2020; Ivanov et al., 2020).
Q2: What GenAI technologies should we focus on in this review?
According to contemporary data science and AI research, we primarily focus on the applications of four types of generative models: Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), Generative Pre-trained Transformers (GPT), and Generative Diffusion Models (Bandi et al., 2023; Eigenschink et al., 2021). Section 2 offers detailed reviews of these models, including their history and rationale, from a data science perspective. We do not focus on the application of deep learning models such as Convolutional Neural Network (CNN), Neural Autoregressive Models, and LSTMs, which have content generation capabilities but are not considered genAI models by nature.
Q3: How GenAI technologies are applied in the smart city sector?
Through our preliminary search conducted in Section 1.2, we did not identify any contemporary research explicitly proposing the integration of GenAI models into urban digital twins to facilitate smart city management, operation, and research. In this regard, this paper aims to outline the scope and strategies for potential GenAI-digital twin integration by investigating the data-driven nature, functionality, and challenges faced by modern urban digital twins. Centered around the challenges of urban digital twins identified in Section 1.2, we will review existing GenAI applications in the broader smart city sector that (1) target key urban subsystems identified by previous studies, and (2) are driven by big urban data. Our literature review in Section 3 will provides answers to this question.
Q4: How GenAI technologies can benefit urban digital twins?
To address this question, a comprehensive review of GenAI applications in various urban sectors is required and is systematically presented in Section 3 of this paper. Urban digital twins, inherently data-driven applications, have emerged alongside advancements in urban computing and informatics. The common ground between GenAI models and urban digital twins lies in their shared capability to generate and handle large volumes and varieties of urban data. Consequently, our literature review in Section 3 focuses on aspects related to data enhancement and generation, simulation, and design. The findings and proposed vision for future opportunities arising from the integration of GenAI and digital twins are discussed in detail in Section 4.
The preliminary answers to our questions are used to formulate the research criteria for identifying relevant studies and to shape the direction and focus of this paper. Given the extensive and interdisciplinary nature of our topic, which involves both research and technical implementation across multiple domains, we justify our approach of using the scoping review methodology. This approach aims to define the fundamental scope of the emerging application areas that bridge GenAI models with urban digital twin technologies.
Figure 1: A summary of GenAI-related publications in major urban sectors using Scopus and IEEE Xplore database (Search conducted on July 20th 2024)
1.4.3 Identifying Relevant Studies
Through a series of standard procedures, in combined with the preliminary answers to our research questions in Section 1.4.2, we devise our strategy for identifying relevant studies as the following:
  1. 1.
    Identify relevant databases: Based on the subject matter, we use Scopus and IEEE Xplore as the primary academic database for conducting our search. The search is conducted based upon keywords and search syntax that matches the paper topics (including article titles, abstracts, and keywords).
  2. 2.
    Define keywords and search syntax for formulating queries: We create two categories of search keywords based on our scopes. These categories describe key concepts in both the GenAI sector and the urban science and smart city sector. For the GenAI category, keywords and search syntax are defined using vocabularies and taxonomies that describe generic GenAI concepts (e.g., ‘Generative AI’, ‘GAI’, ‘GenAI’, ‘Generative Deep Learning’) and specific GenAI models (‘GAN’, ‘VAE’, ‘GPT’, ‘LLM’, ‘Diffusion Model’). We also include various variants, abbreviated forms, and synonyms of these keywords in our search syntax. The urban science category includes keywords that define detailed application areas within individual urban subsystems, as well as generic urban planning tasks. Examples of these keywords include ‘Transportation’, ‘Mobility’, ‘Smart Grid’, ‘Building’, ‘Urban Infrastructure’, and ‘Urban Planning’. The definitions of these keywords are inspired by the urban digital twin application areas identified in Table 1.
  3. 3.
    Define Inclusion and Exclusion Criteria: The inclusion criteria prioritize the review of both peer-reviewed journal articles, conference proceedings, and government technical reports from the last 10 years. For scoping purposes, we will briefly discuss ongoing research published as preprints on SSRN and arXiv to outline potential future research directions, but we will not provide in-depth reviews or discussions of these articles in our review section. We exclude non-peer-reviewed articles, online web pages, studies outside the scope of urban research, and publications older than 10 years. Additionally, we exclude studies that focus on the cybersecurity, building construction, structure design, and agriculture aspects of urban digital twins and IoT applications, which, although relevant to smart city advancements, are outside the scope of urban research.
We perform comprehensive searches based on these criteria in the selected databases. The search results are exported into a CSV file produced by reference management software (Mendeley) for charting purposes. An additional search filter includes limiting the publication language to English.
Figure 2: A summary of GenAI-related publications in major urban sectors using Scopus and IEEE xplore database (Search conducted on July 20th 2024)
1.4.4 Selecting Studies to be Included in the Review
We devised our search keywords based on popular types of GenAI models defined in Section 2, as well as urban application areas summarized in Table 1. These keywords were combined with logical operators and the search templates of Scopus and IEEE Xplore to formulate queries. These keywords are combined with logical operators and the search templates of Scopus and IEEE Xplore to formulate quires. With the specified quires, the initial search results yielded a total of 1,391 papers from two databases, covering 5 different urban application areas. The distribution was as follows: 298 papers in transportation and mobility, 512 in urban energy, 125 in urban water, 184 in building and infrastructure, and 272 in urban planning. After removing duplicates and excluding papers unrelated to urban science (e.g., those focused on cybersecurity and rural environmental issues) by scanning journal types, article types, abstracts, and keywords, the total number of selected papers was reduced to 1,125. The revised distribution was 259 papers in transportation and mobility, 479 in urban energy, 105 in urban water, 81 in building and infrastructure, and 201 in urban planning. These remaining papers will undergo further filtering based on human expert review, assisted by Natural Language Processing (NLP) powered Python scripts and pre-trained Large Language Models (LLMs). The NLP and LLM-based tools used to assist in literature screening extend the methodology developed by Tupayachi et al. (2024) through the LLM-powered generation of ontology and knowledge graphs using scientific literature. The final selected literature is summarized in Figure 1. The article summary for individual urban application areas is provided through Figure 2.
1.4.5 Literature Review Structure
The scoping review approach was chosen for its ability to provide a comprehensive overview of a broad and complex research area, identify key concepts, and map the available evidence (Munn et al., 2018). This method allowed us to include a wide range of study designs and sources, ensuring a thorough exploration of the topic. Our search strategy involved a systematic search of multiple databases, grey literature, and reference lists, adhering to predefined inclusion and exclusion criteria. By using the scoping review method, we aimed to identify research gaps, highlight the diversity of GenAI applications in the urban planning and smart city management sectors, and provide valuable insights for future research directions regarding their integration into urban digital twins.
We will begin by reviewing the five major types of GenAI models used to create our search keywords. This review will cover their rationale, capabilities, and popular applications in computer science, and is presented in Section 2. Following this, we will engage in a comprehensive discussion on GenAI applications in various urban areas, utilizing the screened literature we have collected from two academic databases. Given our focus on urban digital twins, we will further select research articles within individual urban application areas that are centered on the autonomous generation of data, scenarios, 3D city models, and designs for smart city advancement to conduct more detailed review and discussion. The in-depth scenario-based review is presented in Section 3.